# The Algorithmic Gaze: An Audit and Ethnography of the LAION-Aesthetics Predictor Model

### v2

@@ 3.3.1. The LAION-Aesthetics Predictor rates western art more highly than art from the Majority World
As seen in Table 1, the departments with images scored $\geq$ 6.5 are primarily European Paintings, Photographs and Drawing & Prints as well as art from Europe, Asia, and America. On the other hand, African, Native American, Oceanian, Egyptian, Islamic, Ancient West Asian, and Greek & Roman Art. This may be partially because the LAP privileges figurative, two-dimensional art, which would also explain why more abstract art (Modern & Contemporary Art) and physical artifacts (European Sculpture & Decorative Arts) scored lower.
@@
A manual inspection indicates that images rated $\geq$ 6.5 are typically two-dimensional realistic representations of landscapes (e.g., paintings of mountains) or portraits (e.g., photographs or paintings of people). This can also be seen in the artists with multiple images in the top 177. The top two artists with the most images rated $\geq$ 6.5, Katsushika Hokusai (32) and Utagawa Hiroshige (23), are mid-19th century ukiyo-e woodblock print artists who focus on Japanese landscapes (Kafū et al., 2012). Likewise, the artists with the third and sixth most images rated $\geq$ 6.5 are the landscape painters William Trost Richards (6) and Albert Bierstadt (3). The artists with the fourth (Southworth and Hawes, 4) and fifth (Suzuki Shin’ichi, 4) most images in the top 177 are portrait photographers. While the MET dataset contains a broad range of art across cultures and mediums, those that LAP rated $\geq$ 6.5 were mostly paintings or photographs by western or Japanese artists. In the next section, we examine how LAP evaluates paintings in greater detail through the WikiArt dataset.
Median Score % of Pieces Scored $\geq$ 6.5
Genre Cityscape 5.97 539 of 4602 (11.7%)
Portrait 5.90 693 of 14112 (4.9%)
Landscape 5.87 505 of 13358 (3.8%)
Illustration 5.36 11 of 1902 (0.6%)
Abstract Painting 4.94 0 of 4968 (0%)
Style Realism 5.87 634 of 10733 (5.9%)
Romanticism 5.87 371 of 7019 (5.3%)
Impressionism 5.89 646 of 13060 (4.9%)
Cubism 5.50 12 of 2235 (0.5%)
Pop Art 5.05 2 of 1483 (0.1%)
Abstract Expressionism 4.96 0 of 2782 (0%)
Artist Antoine Blanchard 6.94 168 of 170 (98.8%)
Edouard Cortes 7.09 208 of 214 (97.2%)
Pablo Picasso 5.51 1 of 762 (0.1%)
Andy Warhol 5.22 0 of 107 (0%)
Salvador Dalí 5.17 0 of 479 (0%)
Table 2. Selection of highest and lowest scored images across different genres, styles, and artists in WikiArt.
@@ 3.4. Findings: The WikiArt Dataset
The WikiArt dataset is made of 81,444 pieces of visual art from 128 known artists (excluding unknown artists), primarily western artists from the mid-1800s to mid-1900s. Each piece is assigned one of 27 styles across 10 known genres (excluding unknown genres). Genres describe the content of a piece (e.g., landscape or portrait) meanwhile style describes the piece’s form (e.g., realism or impressionism).
@@
As seen in Table 2, we find that LAP highly rates visual art of Cityscapes, Landscapes, and Portraits. These three genres account for 39% of all images in WikiArt, but these genres make up 73% of all images rated $\geq$ 6.5. Likewise, we found that more realistic and representational styles, such as Realism, Romanticism, and Impressionism were highly rated. Here, both style and genre are linked as portraits, landscapes, and cityscape are common subjects in realist, romantic, and impressionistic art. On the other hand, LAP rates less representational art, such as cubism or abstract art, lower. The consequences of this aesthetic bias can be seen across artists. Nearly every piece by the french cityscape artists Edouard Cortès (208 of 214) and Antoine Blanchard (168 of 170) were rated $\geq$ 6.5. In contrast, more influential but less realistic artists—Pablo Picasso, Andy Warhol, and Salvador Dalí— rated lower.
3.5. Findings: Qualitative Visual Analysis
@@ 6.5 $\leq$ score [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
Table 3. Representative sampled images by aesthetic score across the three datasets showing that realist images of landscapes, cityscapes, or people are most highly rated.
@@ 4. How the LAION-Aesthetics Predictor was Made
Seaver (2017) suggests that audits can demonstrate the existence of biases, but “what [audits] cannot do is explain conclusively how that disparate impact came about.” In this section, we explore the potential origins of the aesthetic biases described above by conducting a digital ethnography of how the LAP was made.
4.1. Ethnographic Methodology
Ethnography is a qualitative method from anthropology seeking to explain how groups of people interpret the world (Dourish, 2014). While ethnography began within physical field sites, the 1990s saw a proliferation of multi-sited ethnographic research (Marcus, 1995). Our approach aligns with what Marcus (1995) describes as “following the thing,” a tactic often used to study supply chains (Tsing, 2015). We also draw on the later development of trace ethnography (Geiger and Ribes, 2011), which follows digital documentation, blogs, and code repositories to understand collaboration. Specifically, we draw on: (1) the LAION-Aesthetics announcement blog post (Schuhmann, 2022), (2) the code and data repositories hosting the LAP model, [^3] https://github.com/christophschuhmann/improved-aesthetic-predictor SAC dataset, [^4] https://github.com/JD-P/simulacra-aesthetic-captions AVA dataset, [^5] https://github.com/imfing/ava_downloader and LAION-Logos dataset, [^6] https://huggingface.co/datasets/ChristophSchuhmann/aesthetic-logo-ratings (3) the academic paper for the AVA dataset (Murray et al., 2012), and (4) an  10 minute YouTube video explaining how LAP was made. [^7] https://www.youtube.com/watch?v=OZ3krwyoeWw Our analysis (summarized in Table 4)focused on understanding how LAP and its training data were created and where the images and annotations in LAP’s training data come from.
@@ 4.2.1. Model development choices were made according to the individual taste of LAION’s founder
Schuhmann seems to recognize various limitations of LAP. For example, he encourages others to “experiment” with different model architectures, newer clip embedding models, and larger aesthetic evaluation datasets. Additionally, Schuhmann asks for volunteers to reproduce the training data curation steps described in the video because “[he] had some ugly code somewhere but [he] can’t find it anymore.” He also asks for volunteers in the last comment of the explainer video: “This is all an ongoing area of research and it would be really cool to have someone working on this.” Schuhmann appears to recognize the slapdash way LAP was made and expects that it will be improved upon in the future. However, at the time of our writing in January 2026, LAION has not released a new aesthetic evaluation model.
@@ 4.2.2. Training datasets are inconsistently documented
We initially struggled to find LAION-Logos. While AVA and SAC are linked in the LAION-Aesthetics blog post, LAION-Logos is not. Searching “LAION-Logos” on Google yielded only websites and academic articles alluding to the original blog. In response, we searched the public LAION Discord, eventually finding an unlisted YouTube video posted by Schuhmann in 2022.999https://www.youtube.com/watch?v=OZ3krwyoeWw Because this video is unlisted, it is not searchable on YouTube or indexed by search engines. In the video, Schuhmann explains that LAION-Logos is located on Hugging Face. [^10] https://huggingface.co/datasets/ChristophSchuhmann/aesthetic-logo-ratings LAION-Logos is sparsely documented with a one sentence explanation and rarely used, downloaded only 40 times in the month prior at the time of our writing. While LAION-Logos is technically publicly available, our challenges finding the dateset suggests that it has been de facto private. To the best of our knowledge, we are the first to document LAION-Logos.
@@ 4.2.3. Training datasets were created for different purposes across different time periods
In contrast, AVA was collected a decade prior to popularization of generative AI research. To motivate the construction of AVA, Murray et al. (2012) suggest that “With the ever-expanding volume of visual content available, the ability to organize and navigate such content by aesthetic preference is becoming increasingly important.” The authors position AVA as a solution to the problem of information overload, a common refrain in recommendation research (Seaver, 2022). Perhaps because AVA was created within a personalization paradigm, AVA embodies a more pluralistic conception of aesthetics than the universalism of LAION-Logos and SAC. We describe these competing conceptions of aesthetics below in greater detail.
@@ 4.3.1. Images and aesthetic annotations were collected with mixed levels of consent
Of the 428,632 images used to train LAP, 255,530 ($\sim$60%) come from AVA, 146,372 ($\sim$34%) from SAC, and 26,730 ($\sim$6%) from LAION-Logos. Both the images and ratings from AVA are from photographers who participated in the English-language online art competition website dpchallenge.com between its founding in 2006 and its scraping in 2012. The images in SAC were generated and rated by a handful of AI enthusiasts in 2022 using an early version of Stable Diffusion, with the top 50 most active volunteers accounting for 87.4% of all ratings. Finally, the images in LAION-Logos from LAION-5B were scraped from the internet no later than ealry 2022 and rated by only 18 LAION volunteers. The images used to train LAP have a number of limitations. Temporally, most images are from before 2012, meaning newer content may be underrepresented. Also, the over-representation of photographs in LAP’s training data may explain why the model highly rates photorealism.
@@
Per the LAION-Aesthetics blog, LAP was trained to ”predict the rating people gave when they were asked ’How much do you like this image on a scale from 1 to 10?’” (Schuhmann, 2022). While it is true that SAC and LAION-Logos measures aesthetics as an absolute image property, this is not what AVA measures. Votes on dpchallenge do not measure the perceived aesthetic quality of a de-contextualized image but rather the quality of an image relative to a topic, like skyscapes. As prizes are award to images based on the average final ranking of aesthetic votes for a challenge, votes are comparable within a topic but not between topics. Murray et al. (2012) find that images in challenges with negative affective valences, like fear, tend to receive lower absolute scores. This implies that a model built on absolute scores rather than relative scores within challenges may be confounded by the emotional valence of an image.
@@ 5.1.1. The Imperial Gaze
In our analysis of art from the MET, all of the images rated 6.5+ were paintings, photographs and prints from western or Japanese artists. Not a single piece of African, Oceanian, Native American, Islamic, Egyptian, or West Asian art was rated 6.5+. Although the highly-rated Japanese art may seem like an outlier, western artists have been influenced by a racist fascination with “eastern” (Said, 1979) aesthetics since at least the 1600s (Porter, 1999). Overall, LAP embodies what the post-colonial scholar Kaplan (2012) refers to as the imperial gaze, wherein art is judged according white western tastes. Due to the imperial gaze of LAP, the widespread use of this model to curate training data or evaluate models may exacerbate existing multi-modal biases, like struggling to generate Native American (Taylor et al., 2025) or Persian (Qadri et al., 2025) art.
Similar to prior work on biases in data “quality” filtering (Dodge et al., 2021; Hong et al., 2024), our findings demonstrate that aesthetic evaluation can lead to representational harms by filtering out non-western art. We encourage future research into how other popular image evaluation methods, such as FID scores (Heusel et al., 2017) or CLIP similarity (Radford et al., 2021), asses non-western art. At the same time, we encourage researchers to explore aesthetic biases outside of western ontologies. Notably, none of the images in the MET department of “Africa, the Ancient Americas, and Oceania” scored 6.5+ (Table 1). However, the consolidation of art from different continents and time periods into a monolithic department is a vestige of the colonialist category of “primitive art” (Price, 2001). By using the MET to evaluate LAP, our analysis inherits this coloniality. In calling for a research agenda focused on aesthetic biases, we caution against reinscribing the imperial gaze embedded in western art history.
@@ 5.1.2. The Realist Gaze
So far, we discussed the harms of being excluded by LAP’s algorithmic gaze. However, realism biases can also lead to harm through non-consensual inclusion. We found that the LAION-Aesthetics 6.5+ Dataset is disproportionally made of images from independent visual artists and photographers. Not only have these communities lost jobs due to visual generative AI (Merchant, 2025), but this suggests that their art is driving the development of these technologies. As increasing photorealism exacerbates the greatest harms of visual generative AI—like deepfakes (Mink et al., 2024), job replacement (Jiang et al., 2023), and sexual abuse imagery (Abdulla, 2025)—we encourage researchers to reconsider the value placed on realism in visual evaluation.
@@ 5.1.3. The Male Gaze
We found that images with captions mentioning women were more likely to included in the LAION-Aesthetics 6.5+ Dataset, while those mentioning men or LGBTQ+ people were more likely to be excluded. It follows that datasets curated and models evaluated with LAP will reflect these biases. In doing so, LAP may be reinforcing the heterosexual male gaze in western visual art history (Mulvey, 1975; Berger, 2008) at scale. Although AI fairness researchers often critique data exclusions, Hoffmann (2021) critiques including marginalized communities in oppressive technologies. Recently, women have been immensely harmed by the proliferation of visual generative AI models. The United Nations has warned that generative AI is a major contributor to violence against women (Carmo, 2025). This violence has been facilitated by a rise of commercial AI-based nudification apps, often specifically designed to create nude images of women (Gibson et al., 2025). However, this problem is not unique to dedicated nudification apps. At the time of our writing, thousands of men on X (formerly Twitter) have began using Grok to create and share NCII and CSAM of women and girls (McMahon and Cress, 2026). Aesthetic evaluation models that increase the likelihood of women being included in AI training data, like LAP, risk exacerbating these harms. We caution that greater inclusion in the training data for harmful AI systems can do more harm than good.

### v3

@@ 3.3.1. The LAION-Aesthetics Predictor rates western art more highly than art from the Majority World
As seen in Table 1, the departments with images scored $\geq$ 6.5 are primarily European Paintings, Photographs and Drawing & Prints as well as art from Europe, Asia, and America. On the other hand, not a single piece of African, Native American, Oceanian, Egyptian, Islamic, Ancient West Asian, and Greek & Roman Art was rated 6.5+. This bias is even more apparent when looking at the images rated 6+. Approximately 97% of all images scored 6+ from the MET (3180 of 3284) come from only five departments: (1) Asian Art, (2) Photographs, (3) European Paintings, (4) The American Wing, and (5) Drawing and Prints. This may be partially because the LAP privileges figurative, two-dimensional art, which would also explain why more abstract art (Modern & Contemporary Art) and physical artifacts (European Sculpture & Decorative Arts) scored lower.
@@
A manual inspection indicates that images rated $\geq$ 6.5 are typically two-dimensional realistic representations of landscapes (e.g., paintings of mountains) or portraits (e.g., photographs or paintings of people). This can also be seen in the artists with multiple images in the top 177. The top two artists with the most images rated $\geq$ 6.5, Katsushika Hokusai (32) and Utagawa Hiroshige (23), are mid-19th century ukiyo-e woodblock print artists who focus on Japanese landscapes (Kafū et al., 2012). Likewise, the artists with the third and sixth most images rated $\geq$ 6.5 are the landscape painters William Trost Richards (6) and Albert Bierstadt (3). The artists with the fourth (Southworth and Hawes, 4) and fifth (Suzuki Shin’ichi, 4) most images in the top 177 are portrait photographers. While the MET dataset contains a broad range of art across cultures and mediums, those that LAP rated $\geq$ 6.5 were mostly paintings, prints, or photographs by western or Japanese artists. In the next section, we examine how LAP evaluates paintings in greater detail through the WikiArt dataset.
@@ 3.4. Findings: The WikiArt Dataset
The WikiArt dataset consists of 81,444 pieces of visual art from 128 known artists (excluding unknown artists), primarily western artists from the mid-1800s to mid-1900s. Each piece is assigned one of 27 styles across 10 genres. Genres describe the content of a piece (e.g., landscape or portrait) meanwhile style describes the piece’s form (e.g., realism or impressionism).
@@
As seen in Table 2, we find that LAP highly rates visual art of Cityscapes, Landscapes, and Portraits. These three genres account for 39% of all images in WikiArt, but these genres make up 73% of all images rated 6.5+. Likewise, we found that more realistic and representational styles, such as Realism, Romanticism, and Impressionism were highly rated. Here, both style and genre are linked as portraits, landscapes, and cityscape are common subjects in realist, romantic, and impressionistic art. On the other hand, LAP rates less representational art, such as cubism or abstract art, lower. The consequences of this aesthetic bias can be seen across artists. Nearly every piece by the french cityscape artists Edouard Cortès (208 of 214) and Antoine Blanchard (168 of 170) were rated 6.5+. In contrast, more influential but less realistic artists—Pablo Picasso, Andy Warhol, and Salvador Dalí—were rated lower.
@@ 6.5 $\leq$ score [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
Table 3. A representative selection of images by aesthetic score across the three datasets showing that realist images of landscapes, cityscapes, or people are most highly rated.
3.5. Findings: Qualitative Visual Analysis
@@ 4. How the LAION-Aesthetics Predictor was Made
Seaver (2017) suggests that audits can demonstrate the existence of biases, but “what [audits] cannot do is explain conclusively how that disparate impact came about.” In this section, we explore the potential origins of the aesthetic biases described above by conducting a trace ethnography of how the LAP was made.
4.1. Trace Ethnographic Methodology
Ethnography is a qualitative method from anthropology seeking to explain how groups of people interpret and act in the world (Dourish, 2014). Ethnographic research typically involves extended participant observation and interviews within physical field-sites, ranging from remote villages (Geertz, 1973) to AI research labs (Forsythe, 2001). While ethnography began within physical field sites, the 1990s saw a proliferation of multi-sited ethnographic research due to the rise of globalization and the internet (Marcus, 1995). Within this paradigm, Geiger and Ribes (2011) propose trace ethnography as “a powerful and flexible methodology, able to turn thin documentary traces into ‘thick descriptions’ (Geertz, 1973) of actors and events that are often invisible in today’s distributed, networked environments.” For example, Geiger and Ribes (2011) reconstruct the blocking of a specific Wikipedia vandal by drawing on thin documentary traces of edit history metadata.
We conducted a trace ethnography of the development of the LAION-Aesthetics Predictor by leveraging a variety of digital documents. Specifically, we draw on: (1) the LAION-Aesthetics announcement blog post (Schuhmann, 2022), (2) the code and data repositories hosting the LAP model, [^4] https://github.com/christophschuhmann/improved-aesthetic-predictor SAC dataset, [^5] https://github.com/JD-P/simulacra-aesthetic-captions AVA dataset, [^6] https://github.com/imfing/ava_downloader and LAION-Logos dataset, [^7] https://huggingface.co/datasets/ChristophSchuhmann/aesthetic-logo-ratings (3) the academic paper for the AVA dataset (Murray et al., 2012), and (4) a $\sim$10 minute YouTube video explaining how LAP was made. [^8] https://www.youtube.com/watch?v=OZ3krwyoeWw Our analysis (summarized in Table 4) focused on understanding how LAP and its training data were created and where the images and annotations in LAP’s training data come from.
@@ 4.2.1. Model development choices were made according to the individual taste of LAION’s founder
Schuhmann seems to recognize various limitations of LAP. He encourages others to “experiment” with different model architectures, newer CLIP embedding models, and larger aesthetic evaluation datasets. Additionally, Schuhmann asks for volunteers to reproduce the training data curation steps described in the video because “[he] had some ugly code somewhere but [he] can’t find it anymore.” He also asks for volunteers in the last comment of the explainer video: “This is all an ongoing area of research and it would be really cool to have someone working on this.” Schuhmann appears to recognize the slapdash way LAP was made and expects that it will be improved upon in the future. However, at the time of our writing in January 2026, LAION has not released a new aesthetic evaluation model.
@@ 4.2.2. Training datasets are inconsistently documented
We initially struggled to find LAION-Logos. While AVA and SAC are linked in the LAION-Aesthetics blog post, LAION-Logos is not. Searching “LAION-Logos” on Google yielded only websites and academic articles alluding to the original blog. In response, we searched the public LAION Discord, eventually finding an unlisted YouTube video posted by Schuhmann in 2022.101010https://www.youtube.com/watch?v=OZ3krwyoeWw Because this video is unlisted, it is not searchable on YouTube or indexed by search engines. In the video, Schuhmann explains that LAION-Logos is located on Hugging Face. [^11] https://huggingface.co/datasets/ChristophSchuhmann/aesthetic-logo-ratings LAION-Logos is sparsely documented with a one sentence explanation and rarely used, downloaded only 40 times in the month prior at the time of our writing. While LAION-Logos is technically publicly available, our challenges finding the dateset suggests that it has been de facto private. To the best of our knowledge, we are the first to document LAION-Logos.
@@ 4.2.3. Training datasets were created for different purposes across different time periods
In contrast, AVA was collected a decade prior to the popularization of generative AI research. To motivate the construction of AVA, Murray et al. (2012) suggest that “With the ever-expanding volume of visual content available, the ability to organize and navigate such content by aesthetic preference is becoming increasingly important.” The authors position AVA as a solution to the problem of information overload, a common refrain in recommendation research (Seaver, 2022). Perhaps because AVA was created within a personalization paradigm, AVA embodies a more pluralistic conception of aesthetics than the universalism of LAION-Logos and SAC. We describe these competing conceptions of aesthetics below in greater detail.
@@ 4.3.1. Images and aesthetic annotations were collected with mixed levels of consent
Of the 428,632 images used to train LAP, 255,530 ($\sim$60%) come from AVA, 146,372 ($\sim$34%) from SAC, and 26,730 ($\sim$6%) from LAION-Logos. Both the images and ratings from AVA are from photographers who participated in the English-language online art competition website dpchallenge.com between its founding in 2006 and its scraping in 2012. The images in SAC were generated and rated by a handful of AI enthusiasts in 2022 using an early version of Stable Diffusion, with the top 50 most active volunteers accounting for 87.4% of all ratings. Finally, the images in LAION-Logos from LAION-5B were scraped from the internet no later than early 2022 and rated by only 18 LAION volunteers. The images used to train LAP have a number of limitations. Temporally, most images are from before 2012, meaning newer content may be underrepresented. Also, the over-representation of photographs in LAP’s training data may explain why the model highly rates photorealism.
@@
Per the LAION-Aesthetics blog, LAP was trained to “predict the rating people gave when they were asked ’How much do you like this image on a scale from 1 to 10?”’ (Schuhmann, 2022). While it is true that SAC and LAION-Logos measures aesthetics as an absolute image property, this is not what AVA measures. Votes on dpchallenge do not measure the perceived aesthetic quality of a de-contextualized image but rather the quality of an image relative to a topic, like skyscapes. As prizes are award to images based on the average final ranking of aesthetic votes for a challenge, votes are comparable within a topic but not between topics. Murray et al. (2012) find that images in challenges with negative affective valences, like fear, tend to receive lower absolute scores. This implies that a model built on absolute scores rather than relative scores within challenges may be confounded by the emotional valence of an image.
@@ 5.1.1. The Imperial Gaze
In our analysis of art from the MET, all of the images rated 6.5+ were paintings, photographs and prints from western or Japanese artists. Not a single piece of African, Oceanian, Native American, Islamic, Egyptian, or West Asian art was rated 6.5+. While the highly-rated Japanese art may seem like an outlier, western artists have been influenced by a racist fascination with “eastern” (Said, 1979) aesthetics since at least the 1600s (Porter, 1999). LAP embodies what Kaplan (2012) refers to as the imperial gaze, wherein art is judged according white western tastes. Due to the imperial gaze of LAP, the widespread use of this model to curate training data may exacerbate existing multi-modal biases, like struggling to generate Native American (Taylor et al., 2025) or Persian (Qadri et al., 2025) art.
Similar to prior work on biases in data “quality” filtering (Dodge et al., 2021; Hong et al., 2024), our findings demonstrate that aesthetic evaluations may lead to representational harms by filtering out non-western art. We encourage future research into how other image evaluation methods, such as FID scores (Heusel et al., 2017) or CLIP similarity (Radford et al., 2021), asses non-western art. At the same time, we encourage researchers to explore aesthetic biases outside of western ontologies. Notably, none of the images in the MET department of “Africa, the Ancient Americas, and Oceania” scored 6+ (Table 1). However, the consolidation of art from different continents and time periods into one department is a vestige of the colonialist category of “primitive art” (Price, 2001). By using the MET to evaluate LAP, our analysis inherits this coloniality. In calling for a research agenda focused on aesthetic biases, we caution against reinscribing the imperial gaze embedded in western art history.
@@ 5.1.2. The Realist Gaze
So far, we discussed the harms of being excluded by LAP’s algorithmic gaze. However, realism biases can also lead to harm through non-consensual inclusion. We found that the LAION-Aesthetics 6.5+ Dataset is disproportionally made of images from independent visual artists and photographers. Not only have these communities lost jobs due to visual generative AI (Merchant, 2025), but this suggests that their art may be driving the development of these technologies. As increasing photorealism has the potential to exacerbate the greatest harms of visual generative AI—like deepfakes (Mink et al., 2024), job replacement (Jiang et al., 2023), and sexual abuse imagery (Abdulla, 2025)—we encourage researchers to reconsider the value placed on realism in visual evaluation.
@@ 5.1.3. The Male Gaze
We found that images with captions mentioning women were more likely to included in the LAION-Aesthetics 6.5+ Dataset, while those mentioning men or LGBTQ+ people were more likely to be excluded. It follows that datasets curated and models evaluated with LAP may also reflect these biases. In doing so, LAP reinforces the heterosexual male gaze in western visual art history (Mulvey, 1975; Berger, 2008) at scale. According to Berger (2008) “Women [in western art] are depicted in a quite different way from men — not because the feminine is different from the masculine — but because the ideal spectator is always assumed to be male and the image of the woman is designed to flatter him.” Similarly, we found that LAP was designed according to the taste of one male spectator: LAION’s founder.
The male gaze has long been a feature of computer vision research. In 1973, male researchers at the University of Southern California digitized an image of the Playboy model Lena Forsén to test image compression techniques (Mulvin, 2021). The “Lena image” has since become ubiquitous in computer vision research. Despite decades of criticism that the Lena image perpetuates misogyny in computing (Chang, 2019; Munson, 1996), IEEE publications did not prohibit its use until 2024 (Hern, 2024). Much like Lena’s image circulated without her consent among computer vision researchers for decades, images of (disproportionally) women now similarly circulate among computer vision researchers through the LAION-Aesthetics 6.5+ Dataset.
Although AI fairness researchers often critique data exclusions, Hoffmann (2021) critiques including marginalized communities in oppressive technologies. Recently, women have been immensely harmed by the proliferation of visual generative AI models. The United Nations has warned that generative AI is a major contributor to violence against women (Carmo, 2025). This violence has been facilitated by a rise of commercial AI-based nudification apps, often specifically designed to create nude images of women (Gibson et al., 2025). However, this problem is not unique to dedicated nudification apps. At the time of our writing, thousands of men on X (formerly Twitter) have began using Grok to create and share NCII and CSAM of women and girls (McMahon and Cress, 2026). Of course, we cannot know whether these proprietary models rely on the LAION-Aesthetics Predictor or Datasets. Nevertheless, we expect that aesthetic evaluation models that increase the likelihood of women being included in AI training data risks exacerbating these harms. Much like recent work on how concept filtering impacts T2I CSAM generation (Cretu et al., 2025), we encourage future research on the relationship between aesthetic filtering and the generation of sexual imagery.